OpenStudy (anonymous):

The area of a square is 100. The length of its diagonal is approximately

OpenStudy (anonymous):

12 13 14 15

OpenStudy (anonymous):

area is 100 so side will be 10 length of diagonal will be 10sqrt2

OpenStudy (anonymous):

If the area is 100, the sides are 10. Meaning the diagonal is sqrt(10^2+10^2) = sqrt(200) ~= 14.

OpenStudy (anonymous):

sqrt 2 = 1.414

OpenStudy (anonymous):

so 10 * 1.414 = 14.14 approx = 14

OpenStudy (anonymous):

thanks